a car going around a circular racetrack of radius 100m at 20 m/s2 experiences an acceleration of 4m/s2. what would be the acceleration if the radius of the track was increased to 200m?
1
Expert's answer
2020-06-29T14:03:59-0400

Given, tangential velocity is "v=20m\/s" .

Now, if radius was actually "r=200m" and Tangential acceleration is "a_t=4m\/s^2"

Now, radial acceleration will be,

"a_r=\frac{v^2}{r}"

Thus, "a_r=2m\/s^2" .

Therefore, acceleration of the car will be

"a=\sqrt{a_t^2+a_r^2}=2\sqrt{5}m/s^2"